# A greedy heuristic for graph burning

### v1

@@ Abstract
Given a graph $G$, the optimization version of the graph burning problem seeks for a sequence of vertices, $(u_{1},u_{2},...,u_{k})\in V(G)^{k}$, with minimum $k$ and such that every $v\in V(G)$ has distance at most $k-i$ to some vertex $u_{i}$. The length $k$ of the optimal solution is known as the burning number and is denoted by $b(G)$, an invariant that helps quantify the graph’s vulnerability to contagion. This paper explores the advantages and limitations of an $\mathcal{O}(mn+kn^{2})$ deterministic greedy heuristic for this problem, where $n$ is the graph’s order, $m$ is the graph’s size, and $k$ is a guess on $b(G)$. This heuristic is based on the relationship between the graph burning problem and the clustered maximum coverage problem, and despite having limitations on paths and cycles, it found most of the optimal and best-known solutions of benchmark and synthetic graphs with up to 102400 vertices.
@@ 1 Introduction
The optimization version of the graph burning problem (GBP) is an NP-hard minimization problem that helps quantify a graph’s vulnerability to contagion. It was introduced as a social contagion model [1] and a processor communication problem [2]. Nevertheless, it is an abstract problem with application in many other contexts, particularly those related to information or influence spreading. In detail, GBP receives a graph $G$ as input and seeks a burning sequence of minimum length, where a burning sequence is a list of vertices that burns all the graph’s vertices by the end of the burning process. The latter consists of burning the neighbors of the previously burned vertices and the current vertex in the sequence at discrete time steps. Figure 1 shows the burning process of the optimal burning sequence of a given input graph. The length of the optimal burning sequence of a graph $G$ is known as its burning number and is denoted by $b(G)$.
@@
GBP has an interesting relationship with other NP-hard problems, such as minimum dominating set (MDSP) [3], minimum set cover (MSCP) [4], vertex k-center (VKCP) [5, 6, 7, 8], and firefighter (FP) problems [9]. MDSP can be seen as the GBP with a parallel diffusive process; all vertices are burned in one discrete time step. VKCP reduces to a series of MDSPs and, by transitivity, to a series of MSCPs [5]. Besides, its approximation algorithms are similar to those for GBP [10, 11, 12]. Regarding FP, it aims to protect vertices given the same propagation model used by GBP [9]. This paper shows how GBP is also related to a clustered maximum coverage problem (CMCP), a generalization of the NP-hard maximum coverage problem (MCP) [13, 14]. The remaining part of the paper is organized as follows. Section 2 summarizes the related work, introduces an integer linear program for GBP, and lists the main definitions used in the following sections. Section 3 shows how GBP reduces to a series of CMCPs. Section 4 introduces a $(1/2)$-approximation algorithm for CMCP and a deterministic greedy heuristic for GBP based on this same approximation algorithm. The proposed heuristic is conceptually simple and easy to implement. Besides, it performed unexpectedly well on benchmark and synthetic graphs, as reported in Section 5.
@@ 2 Related work
The burning number conjecture (BNC) is one of the main open questions in the GBP literature. It states that the burning number of any connected graph $G$ is upper bounded by $\lceil n^{1/2}\rceil$ [15], where $n=|V(G)|$. The best-known bound for GBP over arbitrary connected graphs is $\lceil(4n/3)^{1/2}\rceil+1$ [16]. Much research has focused on complexity analysis and finding upper bounds for the burning number on specific graph families, including paths [17], path-forests [17, 11, 18], trees [17], grids [19], intervals [19], fences [20], theta graphs [21], spiders [18], caterpillars [22], graph products [23], generalized Petersen graphs [24], and Jahangir graphs [25]. Parameterized complexity has been studied too [26, 27]. Generalizations of GBP have also been proposed, such as the $w$-graph burning problem ($w$GBP), where $w$ vertices can be selected at each discrete time step [28]; the $r$-graph burning problem ($r$GBP), where a vertex gets burned only if it is adjacent to at least $r$ burned vertices [29]; and the independent cascade graph burning problem (ICGBP), where a vertex $v$ can burned by one of its neighbors $u$ only if the influence $u$ exerts over $v$ is larger than a given threshold [30]. Some heuristics, metaheuristics, approximation algorithms, and mathematical formulations have been proposed to find feasible solutions for GBP. Among the state-of-the-art heuristics and metaheuristics are a centrality-based genetic algorithm (CBAG) [31], a maximum eigenvector centrality heuristic (MECH) [32], a cutting corners heuristic (CCH) [32], a greedy algorithm with forward-looking search strategy (GFSSH) [32], a backbone-based greedy heuristic (BBGH) [33], an improved cutting corners heuristic (ICCH) [33], and a component-based recursive heuristic (CBRH) [33]. Through experimental results, these have proven effective in finding feasible solutions. Among the approximation algorithms are proposals with an approximation ratio of 3 [11] and $3-2/b(G)$ [12] for arbitrary graphs, of 2 for trees [11] and square grids [19], of 1.5 for disjoint paths [11], and polynomial time approximation schemes (PTAS) for linear forests [11] and graphs with small treewidth [34]. Regarding mathematical formulations, an integer linear program (ILP) and two constraint satisfaction problems (CSPs) have been reported [35]. To complement the formal statement of the problem, let us introduce the following ILP, which is equivalent to but simpler than the one previously reported in the literature [35].
min $\ \ \sum\limits_{j\in[1,U]}b_{j}$ (1)
s.t. $\ \ b_{j}\geq 1-b_{ij}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (2)
$\ \ b_{ij}\leq s_{ij}+\sum\limits_{v_{k}\in N[v_{i}]}b_{kj-1}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (3)
$\ \ \sum\limits_{v_{i}\in V}s_{ij}=w$ $\forall j\in[1,U]$ (4)
where $\ \ s_{ij}\ ,\ b_{ij}\ ,\ b_{j}\in\{0,1\}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (5)
$\ \ b_{i0}=0$ $\forall v_{i}\in V$ (6)
@@
Given an input graph $G=(V,E)$, let $n=|V|$, $m=|E|$, and $V=\{v_{1},v_{2},...,v_{n}\}$. This ILP has $\mathcal{O}(Un)$ binary decision variables and $\mathcal{O}(Un)$ constraints, where $U$ is an upper bound on $b(G)$, i.e., $b(G)\leq U\leq n$. Variables $b_{ij}$ (resp., $s_{ij}$) codify which vertices $v_{i}$ are burned (resp., added to the sequence) at each discrete time step $j$. Specifically, $b_{ij}=1$ (resp., $s_{ij}=1$) if vertex $v_{i}$ is burned (resp., is in the sequence) at the discrete time step $j$; $b_{ij}=0$ (resp., $s_{ij}=0$) otherwise. The constant $w$ serves to generalize the program to $w$GBP. If $w=1$, this ILP corresponds to the original GBP. Constraints (2) set $b_{j}=1$ if some vertex is not burned at the discrete time step $j$; since the Objective function (1) is minimized, if all vertices are burned at the discrete time step $j$, then $b_{j}=0$. This way, the objective function seeks to burn all vertices as soon as possible, i.e., at the smallest index $j$. Constraints (3) guarantee that a vertex cannot be burned unless it is into the sequence, was already burned, or has some burned neighbor. Finally, Constraints (4) indicate that $w$ vertices must be added into the sequence at each discrete time step. The solution is codified into variables $s_{ij}$ and $b(G)$ equals the smallest index $j$ at which all vertices are burned; namely, $b(G)$ is the value of the objective function plus one. Before finishing this section, let us list the basic definitions used in the remaining part of the paper.
Basic definitions
@@
The distance between two vertices in a graph is the length of their shortest path.
@@
A burning sequence, of a graph $G=(V,E)$, is a function $f:\{x\in\mathbb{Z}_{+}:x\leq k\}\rightarrow V$ such that Eq. (7) holds. The length of the burning sequence is $k$.
$\bigcup_{i=1}^{k}N_{k-i}[f(i)]=V$ (7)
@@
Given an input graph, GBP seeks a burning sequence of minimum length.
@@ 3 Burning and covering
The clustered maximum coverage problem (CMCP) receives $k$ clusters as input, each containing subsets of a universe of elements. Its goal is to select one subset from each cluster so that the cardinality of their union is maximized.
@@
MCP receives a collection of subsets and a positive integer $k$ as input. Its goal is to find at most $k$ subsets such that the cardinality of their union is maximized [13]. Notice that an instance for MCP can be transformed into one for CMCP by creating $k$ identical copies of the input collection. Under these conditions, solving CMCP is equivalent to solving MCP. ∎
Since MCP is NP-hard, CMCP is too. The best approximation ratio achievable for the former cannot be greater than $(1-\frac{1}{e})$ unless $P=NP$ [13, 36, 14]. Thus, by Proposition 1, CMCP has the same inapproximability results.
@@
Let $V$ be the universe of elements for CMCP, where $G=(V,E)$ is the input graph for GBP. The clusters for the corresponding CMCP are $C_{r}=\{N_{r}[v]:v\in V\},\forall r\in[0,k-1]$, where $k$ is a guess on $b(G)$; breadth-first search [37, 38] can be used to compute the distance between all pairs of vertices efficiently. An optimal solution for CMCP consists of one subset from each cluster. An optimal solution for GBP is obtained by properly concatenating the corresponding vertices of each selected subset (See Definition 8). For instance, the clusters for a path $P_{4}=(\{v_{1},v_{2},v_{3},v_{4}\},\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\}\})$, with $k=b(P_{4})=2$, are $C_{0}=\{N_{0}[v_{1}],N_{0}[v_{2}],N_{0}[v_{3}],N_{0}[v_{4}]\}$ and $C_{1}=\{N_{1}[v_{1}],N_{1}[v_{2}],N_{1}[v_{3}],N_{1}[v_{4}]\}$. An optimal solution for this instance is $\{N_{0}[v_{4}],N_{1}[v_{2}]\}$, which covers all the vertices, and its corresponding optimal burning sequence is $(v_{2},v_{4})$. Naturally, solving CMCP with $k=b(G)$ is necessary for the reduction to be useful. To find that value, we can solve CMCP within a binary search between 1 and $n$ (See Algorithm 1). If $k\geq b(G)$, an optimal solution for CMCP must cover all the vertices (line 8). If the optimal solution for CMCP does not cover all vertices, then $k<b(G)$ (line 10). This way, $k=b(G)$ at some point in the binary search, and the optimal burning sequence is found. ∎
@@
7:     $k\leftarrow\lfloor(l+h)/2\rfloor$
8:     $s\leftarrow$ ExactSolver$(\text{CMCP}(G,k,d))$
@@
To further explain this reduction, let us introduce an ILP for GBP as a CMCP (See Expressions (8) to (11)). The clusters are constructed based on the input graph $G=(V,E)$ and a given integer $k$; let us assume that $V=\{v_{1},v_{2},...,v_{n}\}$. This program has $\mathcal{O}(kn)$ binary decision variables and $\mathcal{O}(n)$ constraints. Decision binary variables $x_{rj}$ are used to indicate which subset from each cluster is part of the solution (See Constraints (9)). Actually, this is an ILP for $w$GBP; nevertheless, in this paper, we stick to the original GBP, i.e., $w=1$. Decision binary variables $b_{j}$ indicate which vertices the selected subsets cover. Since the Objective function (8) maximizes the sum over all $b_{j}$, we add Constraints (10) to avoid a vertex $v_{j}$ being marked as covered (i.e., $b_{j}=1$) if it is not an element of some of the selected subsets; notice that |$|\{N_{r}[v]:v\in V\}|=n$ for each cluster $r\in[0,k-1]$. Of course, a burning sequence of length $k$ exists only if $k\geq b(G)$. If so, the burning sequence is codified into variables $x_{ri}$. Explicitly, for $w=1$, the burning sequence is $(v_{g(x_{k-1})},v_{g(x_{k-2})},...,v_{g(x_{0})})$, where $g(x_{r})$ is the index $i$ where $x_{ri}=1$. In case $k<b(G)$, a burning sequence of such length does not exist.
@@
s.t. $\ \ \sum\limits_{v_{i}\in V}x_{ri}=w$ $\forall r\in[0,k-1]$ (9)
$\ \ b_{j}\leq\sum\limits_{r\in[0,k-1]}\ \ \sum_{v_{i}\in N_{r}[v_{j}]}x_{ri}$ $\forall v_{j}\in V$ (10)
where $\ \ x_{rj}\ ,\ b_{j}\in\{0,1\}$ $\forall r\in[0,k-1],\forall v_{j}\in V$ (11)
The next section introduces a greedy $(1/2)$-approximation algorithm for CMCP and its heuristic adaptation to GBP.

### v3

@@ Abstract
Given a graph $G$, the optimization version of the graph burning problem seeks for a sequence of vertices, $(u_{1},u_{2},...,u_{p})\in V(G)^{p}$, with minimum $p$ and such that every $v\in V(G)$ has distance at most $p-i$ to some vertex $u_{i}$. The length $p$ of the optimal solution is known as the burning number and is denoted by $b(G)$, an invariant that helps quantify the graph’s vulnerability to contagion. This paper explores the advantages and limitations of an $\mathcal{O}(mn+pn^{2})$ deterministic greedy heuristic for this problem, where $n$ is the graph’s order, $m$ is the graph’s size, and $p$ is a guess on $b(G)$. This heuristic is based on the relationship between the graph burning problem and the clustered maximum coverage problem, and despite having limitations on paths and cycles, it found most of the optimal and best-known solutions of benchmark and synthetic graphs with up to 102400 vertices. Beyond practical advantages, our work unveils some of the fundamental aspects of graph burning: its relationship with a generalization of a classical coverage problem and compact integer programs. With this knowledge, better algorithms might be designed in the future.
@@ 1 Introduction
The optimization version of the graph burning problem (GBP) is an NP-hard minimization problem that helps quantify a graph’s vulnerability to contagion. It was introduced as a social contagion model by Bonato et al. [1] and as a processor communication problem by Alon [2]. Nevertheless, it is an abstract problem with application in many other contexts, particularly those related to information or influence spreading [3, 4, 5]. In detail, GBP receives a graph $G$ as input and seeks a burning sequence of minimum length, where a burning sequence is a list of vertices that burns all the graph’s vertices by the end of the burning process. This process consists of burning the neighbors of the previously burned vertices and the current vertex in the sequence at discrete time steps. Figure 1 shows the burning process of the optimal burning sequence of a given input graph. The length of the optimal burning sequence of a graph $G$ is known as its burning number and is denoted by $b(G)$.
@@
The remaining part of the paper is organized as follows. Section 2 summarizes the related work, introduces an integer linear program (ILP) for GBP, and lists the main definitions used in the following sections. Section 3 shows how GBP reduces to a series of clustered maximum coverage problems (CMCP). Section 4 introduces two more ILPs, a $(1/2)$-approximation algorithm for CMCP, and a deterministic greedy heuristic for GBP based on the approximation algorithm for CMCP. The proposed heuristic is conceptually simple and easy to implement. Besides, it performed unexpectedly well on benchmark and synthetic graphs, as reported in Section 5.
@@ 2 Related work
Many of the theoretical and practical aspects of GBP have been deeply explored in the last years. These include complexity analysis on specific graph classes, seeking tighter upper bounds for the burning number, mathematical modeling, and algorithm design. For a comprehensive survey, the reader can refer to the paper of Bonato [6].
GBP has an interesting relationship with other NP-hard problems. For instance, the minimum dominating set problem can be seen as the GBP with a parallel propagation process, where all vertices are burned in one discrete time step [7]. Besides, the known approximation algorithms for GBP are very similar to those for the vertex k-center problem [8, 9, 10], which reduces to a series of minimum dominating set problems and, by transitivity, to a series of minimum set cover problems [11, 12]. This paper shows how GBP is also related to the CMCP, a generalization of the NP-hard maximum coverage problem (MCP) [13, 14].
Regarding heuristics for the GBP, most of them rely on exploiting centrality measures, local search, or evolutionary computing. Therefore, they are relatively complicated and can be very time consuming. Among these are a centrality based genetic algorithm [15] and different heuristics based on eigenvector centrality [16, 3]. In constrast to these heuristics, the one introduced in this paper is simpler, is supported by deductive proofs, and has a good empirical performance, i.e., it found most of the optimal solutions of benchmark datasets in reasonable amounts of time.
Regarding mathematical models for the GBP, an integer linear program (ILP) and two constraint satisfaction problems (CSPs) have been reported [17]. In this paper, we introduce three novel ILPs for GBP with a much better performance when implemented on a commercial optimization solver. Expressions (1) to (6) show the first of these ILPs, named ILP-PROP because it is based on the propagation process of the GBP. In this model, $G=(V,E)$ is the input graph, $n=|V|$, $m=|E|$, and $V=\{v_{1},v_{2},...,v_{n}\}$. It has $2Un+U$ binary decision variables and $2Un+U$ constraints, where $U$ is an upper bound on $b(G)$. Variables $b_{i,j}$ (resp., $s_{i,j}$) codify which vertices $v_{i}$ are burned (resp., added to the sequence) at each discrete time step $j$. Specifically, $b_{i,j}=1$ (resp., $s_{i,j}=1$) if vertex $v_{i}$ is burned (resp., is in the sequence) at the discrete time step $j$; $b_{i,j}=0$ (resp., $s_{i,j}=0$) otherwise. If all vertices are burned by step $j$, $x_{j}=0$; otherwise, $x_{j}=1$.
min $\ \ \sum\limits_{j\in[1,U]}x_{j}+1$ (1)
s.t. $\ \ x_{j}\geq 1-b_{i,j}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (2)
$\ \ b_{i,j}\leq s_{i,j}+\sum\limits_{v_{k}\in N[v_{i}]}b_{k,j-1}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (3)
$\ \ \sum\limits_{v_{i}\in V}s_{i,j}=1$ $\forall j\in[1,U]$ (4)
where $\ \ s_{i,j}\ ,\ b_{i,j}\ ,\ x_{j}\in\{0,1\}$ $\forall v_{i}\in V,\forall j\in[1,U]$ (5)
$\ \ b_{i,0}=0$ $\forall v_{i}\in V$ (6)
@@
Constraints (2) set $x_{j}=1$ if some vertex is not burned at the discrete time step $j$; since the Objective function (1) is minimized, $x_{j}$ is set to zero when all vertices are burned at the discrete time step $j$. This way, the objective function seeks to burn all vertices as soon as possible, i.e., at the smallest index $j$. Constraints (3) guarantee that a vertex cannot be burned unless it is into the sequence, was already burned, or has some burned neighbor. Finally, Constraints (4) indicate that 1 vertex must be added into the sequence at each discrete time step. The solution is codified into variables $s_{i,j}$, and $b(G)$ equals the smallest index $j$ at which all vertices are burned; namely, $b(G)$ is the value of the objective function; notice that the extra 1 in the objective function comes from the fact the smallest index $j$ at which all vertices get burned has $x_{j}=0$.
Before finishing this section, let us list the basic definitions used in the remaining part of the paper.
@@
The distance between two vertices in a graph, $d(u,v)$, is the length of their shortest path.
@@
A burning sequence of a graph $G=(V,E)$ is an ordered list $(u_{1},u_{2},...,u_{p})\in V^{p}$ such that the distance from every $v\in V$ to some vertex $u_{i}$ is at most $p-i$. The length of the burning sequence is $p$.
@@
Given an input graph, the GBP seeks a burning sequence of minimum length.
@@ 3 Burning and covering
The clustered maximum coverage problem (CMCP) receives $p$ clusters as input, each containing subsets of a universe of elements. Its goal is to select one subset from each cluster so that the cardinality of their union is maximized.
@@
MCP receives a collection of subsets and a positive integer $p$ as input. Its goal is to find at most $p$ subsets such that the cardinality of their union is maximized [13]. Notice that an instance for MCP can be transformed into one for CMCP by creating $p$ identical copies of the input collection. Thus, solving the resulting CMCP implies solving MCP. ∎
Since MCP is NP-hard, CMCP is NP-hard too. The best approximation ratio achievable for the former cannot be greater than $(1-\frac{1}{e})$ unless $P=NP$ [13, 18, 14]. Thus, by Proposition 1, CMCP has the same inapproximability results.
@@
Let $V$ be the universe of elements for CMCP, where $G=(V,E)$ is the input graph for GBP. The clusters for the corresponding CMCP are $\mathcal{C}_{r-1}=\{N_{r-1}[v]:v\in V\},\forall r\in[1,p]$, where $p$ is a guess on $b(G)$; breadth-first search [19, 20] can be used to compute the distance between all pairs of vertices efficiently. An optimal solution for CMCP consists of one subset from each cluster. An optimal solution for GBP is obtained by properly concatenating the corresponding vertices of each selected subset (See Definition 8). For instance, the clusters for a path $P_{4}=(\{v_{1},v_{2},v_{3},v_{4}\},\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\}\})$, with $p=b(P_{4})=2$, are $\mathcal{C}_{0}=\{N_{0}[v_{1}],N_{0}[v_{2}],N_{0}[v_{3}],N_{0}[v_{4}]\}$ and $\mathcal{C}_{1}=\{N_{1}[v_{1}],N_{1}[v_{2}],N_{1}[v_{3}],N_{1}[v_{4}]\}$. An optimal solution for this instance is $\{N_{0}[v_{4}],N_{1}[v_{2}]\}$, which covers all the vertices, and its corresponding optimal burning sequence is $(v_{2},v_{4})$. Naturally, solving CMCP with $p=b(G)$ is necessary for the reduction to be useful. To find that value, we can solve CMCP within a binary search between 1 and $n$ (See Algorithm 1). If $p\geq b(G)$, an optimal solution for CMCP must cover all the vertices (line 8). If the optimal solution for CMCP does not cover all vertices, then $p<b(G)$ (line 10). This way, $p=b(G)$ at some point in the binary search, and the optimal burning sequence is found. ∎
@@
7:     $p\leftarrow\lfloor(l+h)/2\rfloor$
8:     $S\leftarrow$ Solve CMCP$(G,p,D)$
@@
To further explain this reduction, let us introduce an ILP for GBP as a CMCP (See Expressions (7) to (10)); we refer to this ILP as ILP-CMCP. The clusters are constructed based on the input graph $G=(V,E)$ and a given integer $p$; let us assume that $V=\{v_{1},v_{2},...,v_{n}\}$. This program has $pn+n$ binary decision variables and $p+n$ constraints. Decision binary variables $x_{i,j}$ are used to indicate which subset from each cluster is part of the solution (See Constraints (8)). Decision binary variables $b_{j}$ indicate which vertices the selected subsets cover. Since the Objective function (7) maximizes the sum over all $b_{j}$, we add Constraints (9) to avoid a vertex $v_{j}$ being marked as covered (i.e., $b_{j}=1$) if it is not an element of some of the selected subsets; notice that |$|\{N_{i-1}[v]:v\in V\}|=n$ for each cluster $i\in[1,p]$. Of course, a burning sequence of length $p$ exists only if $p\geq b(G)$. If so, the burning sequence is codified into variables $x_{i,j}$. Explicitly, the burning sequence is $(v_{g(p)},v_{g(p-1)},...,v_{g(1)})$, where $g(i)$ is the index $j$ where $x_{i,j}=1$. In case $p<b(G)$, a burning sequence of such length does not exist.
@@
s.t. $\ \ \sum\limits_{v_{j}\in V}x_{i,j}=1$ $\forall i\in[1,p]$ (8)
$\ \ b_{j}\leq\sum\limits_{i\in[1,p]}\ \ \sum_{v_{k}\in N_{i-1}[v_{j}]}x_{i,k}$ $\forall v_{j}\in V$ (9)
where $\ \ x_{i,j}\ ,\ b_{j}\in\{0,1\}$ $\forall i\in[1,p],\forall v_{j}\in V$ (10)
ILP-CMCP aims at covering as many vertices as possible and requires knowing $b(G)$ in advance to be equivalent to the GBP. As already mentioned, such an issue can be solved by using a binary search. However, ILP-CMCP can also be modified to remove such requirements. We refer to the resulting ILP as ILP-COV because it is based on the coverage definition of the GBP (See Expressions (11) to (17)).
min $\ \ \sum\limits_{i\in[1,U]}\sum\limits_{v_{j}\in V}x_{i,j}$ (11)
s.t. $\ \ \sum_{v_{j}\in V}x_{i,j}\leq\sum_{v_{j}\in V}x_{i-1,j}$ $\forall i\in[1,U]$ (12)
$\ \ \sum\limits_{v_{j}\in V}x_{i,j}\leq 1$ $\forall i\in[1,U]$ (13)
$\ \ b_{j}\leq\sum\limits_{i\in[1,U]}\ \ \sum_{v_{k}\in N_{i-1}[v_{j}]}x_{i,k}$ $\forall v_{j}\in V$ (14)
$\ \ \sum_{v_{j}\in V}b_{j}=n$ (15)
where $\ \ x_{i,j}\ ,\ b_{j}\in\{0,1\}$ $\forall i\in[1,U],\forall v_{j}\in V$ (16)
$\ \ \sum_{v_{j}\in V}x_{0,j}=1$ (17)
ILP-COV has the same binary variables defined for ILP-CMCP. However, instead of a guess $p$, it requires an upper bound $U$ on $b(G)$. The Objective function (11) minimizes the number of vertices in the burning sequence. Thanks to Constraints (12), (13), and (17), only 1 vertex for each valid coverage in $\{0,1,\dots,b(G)-1\}$ can be selected. By Constraints (14), a vertex cannot be marked as covered if it is not in the $i-1$ neighborhood of some vertex with coverage capacity $i-1$. Finally, Constraint (15) guarantees that all vertices are burned. ILP-COV has $Un+n$ binary variables and $12U+n+$ 1 constraints.
The next section introduces a greedy $(1/2)$-approximation algorithm for CMCP and its heuristic adaptation to GBP. This heuristic was tested in Section 5, and the solutions it returned were used as upper bound for ILP-COV, which was implemented in a commercial optimization solver to find optimal solutions. This way, we could evaluate the quality of the solutions returned by the proposed greedy algorithm.